The internal energy of a gas is given by U = 1.5 pV. It expands from 100 cm3 to 200 cm3 against a constant pressure of 1.0 × 105 Pa. Calculate the heat absorbed by the gas in the process.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. U = 1.5 pV P = 1 × 10^5 Pa dV = (200 – 100) cm^3 = 100 cm^3 = 10^–4 m^3 dU = 1.5 × 10^5 × 10^–4 = 15 dW = 1.5 x 10^5 x 10^-4 = 15 dW = 10^5 x 10^-4 = 10 dQ = dU + dW = 10 + 15 = 25 J